# DBCopilot: Scaling Natural Language Querying to Massive Databases

### v2

@@ A.1. Case Study
DPR $\langle\text{world},\{\text{countrylanguage}\}\rangle$
[…]
Figure 7. Successful case where DBCopilot routes the correct schema but baselines fail.
[…]
We provide successful cases to illustrate how DBCopilot improves schema routing through joint retrieval. As the example in Figure 7 shows, DBCopilot is the only approach that accurately routes the appropriate schema for the given natural language question. Sparse retrieval methods (BM25 and CRUSHBM25) rely heavily on lexical matching while ignoring the semantic matching between the question and the schemas. However, untrained dense retrieval methods (SXFMR and CRUSHSXFMR) also fail to capture the underlying semantics of tables and retrieve the wrong databases. Fine-tuning based approaches identify the correct database where DPR misses the country table linked to the countrylanguage table, but relation-aware DBCopilot does not.
[…]
A.2. Efficiency & Resource Consumption
We also measure the efficiency and resource consumption of each schema routing method. Table 7 shows the number of queries per second (QPS), the duration of build time, the amount of disk space utilized, and the consumption of system and GPU memory. LLM inference and complex post-processing are the main reasons for the slowness of CRUSH approaches. DBCopilot, based on an encoder-decoder architecture, consumes more disk space and GPU memory than dense retrieval approaches (SXFMR, DPR) based on an encoder-only architecture. Despite the slightly higher resource consumption for training and inference in generative retrieval, we believe and have observed that the growing popularity of generative AI is leading to an increasing number of optimizations being proposed.
Table 7. Efficiency & resource consumption of different methods. Build time includes the time for model training and index construction. We calculate queries per second (QPS) and memory consumption with a query batch size of 64.
Method QPS Build (s) Disk (MB) RAM (MB) GPU (MB)
BM25 1677.1 10 36 800 -
SXFMR 49.0 48 420 2,230 3742
CRUSHBM25 0.9 10 36 1104 -
CRUSHSXFMR 0.8 48 420 2502 3742
BM25 1677.1 6972 36 800 -
DPR 49.0 5161 420 2,230 3742
DBCopilot 43.0 13678 852 656 11807
A.3. Oracle Test with In-Context Learning
Table 8. Schema-Aware NL2SQL with In-Context Learning
Provided Schema Prompt Strategy Spider Bird Spidersyn
EX Cost EX Cost EX Cost
G. Column Basic 77.95 3.52 35.98 11.58 76.69 3.43
G. Table Basic 75.82 3.61 28.49 12.10 69.34 3.52
G. Database Basic 75.73 3.83 25.75 13.08 63.64 3.74
5 Database Basic 68.38 5.57 21.45 17.79 50.48 5.47
Table 8 shows the results of the oracle test by providing few-shot (5-shot) examples. As we can see, while this improves the accuracy of NL2SQL somewhat, it does not compensate for the performance degradation caused by extraneous schema elements. This makes schema linking still important in LLM-based NL2SQL. In the future, we will investigate how to better disambiguate and link schema elements in the context of schema-agnostic NL questions.

### v3

@@ 4.4. Ablation Study (RQ5)
We perform ablation studies on DBCopilot to validate the contribution of each module, as summarized in § 4.4.
DFS Serialization.  DFS serialization is essential for schema routers to learn the underlying relations of databases and tables. To demonstrate this, we replace the DFS serialization with the unordered basic serialization of tables, similar to DSI for multiple documents. As shown in § 4.4, this results in a 1% to 6% reduction in recall. In addition, DFS serialization facilitates graph-based constrained decoding to reduce the search space, either by first generating a database and then generating its tables, or by further generating based on existing tables.
Data Synthesis.  Generative retrieval enables relation-aware end-to-end schema routing, but is unable to generalize to unseen schemata during training. We can see that routers trained on instances ($N$, $S$) extracted from original NL2SQL datasets are almost completely incapable of routing on unseen test schemata. Data synthesis by sampling schemata and generating pseudo-questions can effectively inject semantic mappings to model parameters. Besides, mixing original and synthetic data for training does not lead to more accurate schema routing, possibly due to label imbalance. Figure 10 shows how the performance varies with the amount of synthetic data. Database and table recalls increase rapidly at first as more data is synthesized, but level off after a certain amount. How to improve the quality of synthetic data and reduce the need for its quantity may be a future direction.
Decoding Strategy.  Decoding optimizations (i.e., constrained decoding and group beam search) are beneficial for generative schema routing during inference. Specifically, constrained decoding is useful for generating accurate table identifiers, since removing it mainly affects table recalls, about 2% to 4%. Meanwhile, group beam search helps to generate multiple diverse schemata because database recall@5 decreases significantly without it.
Finding 7. DFS serialization, data synthesis, and decoding optimizations contribute in different ways.
5. Related Work
5.1. Natural Language to SQL
NL2SQL is an essential part of natural language interfaces to databases, enabling non-experts to easily query structured data. Most research has approached this task along the lines of semantic parsing (Kamath and Das, 2019; Li et al., 2020). Early NL2SQL parsing systems adopted symbolic approaches that relied heavily on rule or grammar engineering, making them difficult to apply widely (Zelle and Mooney, 1996; Woods, 1973). In recent years, with the advances in deep learning, especially PLMs, NL2SQL approaches that using neural generation models have become the mainstream (Li et al., 2023c; Wang et al., 2020; Gu et al., 2023a; Li et al., 2023a; Yu et al., 2021). Some efforts have been made to improve encoding and decoding, such as relation-aware encoding (Wang et al., 2020; Li et al., 2023a) and grammar-based decoding (Xu et al., 2017; Scholak et al., 2021). Some other efforts have focused on generating accurate SQL queries through schema linking (Li et al., 2023c; Wang et al., 2020) or skeleton generation (Gu et al., 2023a; Gu et al., 2023b).
LLMs have emerged as a new paradigm for NL2SQL (Rajkumar et al., 2022; Liu et al., 2023; Gao et al., 2024; Li et al., 2024a). Unlike previous work, a core component of the LLM-based NL2SQL solution is prompt engineering, i.e., prompting LLM to generate correct and efficient SQL queries. Several efforts have been made to select appropriate in-context learning examples through question classification (Pourreza and Rafiei, 2023), skeleton retrieval (Guo et al., 2023), and in-domain example synthesis (Chang and Fosler-Lussier, 2023). Research has also been done on decoupling zero-shot NL2SQL into PLM-based sketch generation and LLM-based SQL completion (Gu et al., 2023b). Recently, pipeline-based approaches decompose NL2SQL into steps including schema linking, query classification, query decomposition, SQL refinement, further extending the power of LLM-based NL2SQL (Pourreza and Rafiei, 2023; Li et al., 2024b). These studies are orthogonal to our work and can be easily integrated to optimize SQL generation of DBCopilot.
5.2. Table Discovery
Table discovery techniques are critical to data management, allowing users to explore and gain insights from data warehouses or lakes (Fan et al., 2023a). A common practice for table discovery is query-driven discovery (Zhang and Ives, 2020; Nargesian et al., 2019), where the query is typically in the form of keywords, tables, or NL questions. For keyword search, data discovery systems find relevant tables by searching for topic terms in the metadata (Cafarella et al., 2009; Pimplikar and Sarawagi, 2012). Table understanding (Deng et al., 2020), including domain discovery (Li et al., 2017; Ota et al., 2020) and table annotation (Limaye et al., 2010; Venetis et al., 2011), has emerged as a means to mitigate the missing, incomplete, and inconsistent table schemata for keyword search. For table search, tables are retrieved to augment a given table as the query (Sarma et al., 2012). Joinable table search (Dong et al., 2021; Zhu et al., 2019) and unionable table search (Fan et al., 2023b; Khatiwada et al., 2023) aim to augment a query table with additional attributes and new tuples, respectively. For NL question search, open domain question answering retrieves independent tables to answer the NL question with table-based models (Chen et al., 2021; Herzig et al., 2021). Although dense retrieval methods based on contrastive learning have achieved advanced performance in the above tasks (Dong et al., 2021; Fan et al., 2023b; Herzig et al., 2021), they are difficult to consider the structural relationships between tables and retrieve them jointly.
Alternatively, data navigation allows users to navigate interactively through the tables organized hierarchically (Nargesian et al., 2020) or relationally (Fernandez et al., 2018). Our work can be seen as automated data navigation, where we employ relation-aware DSI to model the query-based navigation in an end-to-end manner. There is another line of research that focuses on table content discovery for Web table QA (Badaro et al., 2023), which typically employs PLMs to manipulate and analyze HTML tables of limited size (with average #row < 20) (Hulsebos et al., 2023). However, these systems are not suitable for NLIDBs as they cannot efficiently handle large relational tables with thousands or millions of rows and diverse data types that are better handled through semantic parsing.
6. Discussion & Future Work
DBCopilot has made significant progress in schema routing and schema-agnostic NL2SQL. However, there are some limitations that need to be addressed in future work.
  1. 1)
    Cross-Databases Querying. While current system shows the capability to route NL questions over massive databases, it is limited to generating SQL queries against a single database. Extending DBCopilot to support cross-database SQL querying faces several challenges: i) the heterogeneity of different databases and the complexity of ETL processes for cross-database SQL querying; ii) the increased difficulty in discovering implicit relationships between tables across multiple databases than within a single database; iii) the absence of comprehensive benchmarks for evaluating cross-database NL2SQL systems. We envision an alternative direction to decompose complex NL questions into multiple sub-questions, each answerable by a single database, and then integrate the results.
  2. 2)
    Synthetic Data Quality. As discussed in the case study, the quality of synthetic data used for training has a notable impact on router performance. Future efforts could focus on: i) generating high-quality synthetic data by using larger LLMs with better factual consistency; ii) providing more schema descriptions to reduce hallucination; iii) implementing automated validation mechanisms using TNLI (Bussotti et al., 2023) to identify and remove low-quality synthetic data. In addition, reducing the amount of data used to train schema routers while maintaining their performance is also a worthwhile exploration.
  3. 3)
    Dynamic Schema Update. In real-world scenarios, database schemata may evolve over time, which creates a need for all retrieval-based methods to update the index. Although we have decoupled the dynamic schema routing into a compact and flexible copilot model, and it currently takes less than 4 hours to train a schema router from scratch on thousands of tables, managing to incremental update DSI can further reduce the computational resources required for evolving schemata.
7. Conclusion
In this paper, we present DBCopilot, a framework for effectively scaling natural language querying to massive databases. DBCopilot decouples schema-agnostic NL2SQL into schema routing and SQL generation through LLM-Copilot collaboration. We present an approach to effectively navigate the target database and tables for an NL question in a relation-aware, end-to-end manner. We also propose a training data synthesis paradigm to adapt routers over massive databases without manual intervention, and explore various prompt strategies to incorporate multiple candidate schemata in LLM-based SQL generation. Extensive experiments verify the effectiveness of our solution. In summary, DBCopilot pushes the boundaries of NL2SQL, enabling researchers to better strategize their quest for data accessibility.
Acknowledgements.
We sincerely thank the reviewers for their insightful comments and valuable suggestions. This work was supported in part by Beijing Natural Science Foundation under Grant L243006, and in part by the Natural Science Foundation of China under Grant No. 62476265 and 62306303.